OpenStudy (anonymous):

ln(e^18x) = 9 Solve the equation for x. Round your answer to two decimal places. Do not include "x =" in your answer.

myininaya (myininaya):

\[\ln(e^{18x})=9\] is this right?

OpenStudy (anonymous):

thats the equation, yes

myininaya (myininaya):

if so we have 18x=9 x=2

OpenStudy (anonymous):

its 0.5, thanks for ur help tho

myininaya (myininaya):

oh yes it is lol

OpenStudy (anonymous):

lol

myininaya (myininaya):

18x=9 x=9/18

OpenStudy (anonymous):

24 * log(3x) = 60

myininaya (myininaya):

x=1/2

OpenStudy (anonymous):

can u do that one? its the same thing. solve for x

OpenStudy (anonymous):

i think its 105.41 but im not sure

myininaya (myininaya):

\[\log(3x)=\frac{60}{24}\] \[\log(3x)=\frac{5}{2}\] \[10^{\log(3x)}=10^\frac{5}{2}\] \[3x=10^\frac{5}{2}\] \[x=\frac{1}{3}*10^\frac{5}{2}\]

OpenStudy (anonymous):

YAY! i was right :)

myininaya (myininaya):

\[x=\frac{1}{3}*10^\frac{4}{2}*10^\frac{1}{2}=\frac{1}{3}*10^2*\sqrt{10}=\frac{100}{3}\sqrt{10}\]

OpenStudy (anonymous):

Solution: x = e^8\/18